### v1

# Multivariate Causal Effects: a Bayesian Causal Regression Factor Model

Abstract.
The impact of wildfire smoke on air quality is a growing concern, contributing to air pollution through a complex mixture of chemical species with important implications for public health. While previous studies have primarily focused on its association with total particulate matter (pm2.5), the causal relationship between wildfire smoke and the chemical composition of pm2.5 remains largely unexplored. Exposure to these chemical mixtures plays a critical role in shaping public health, yet capturing their relationships requires advanced statistical methods capable of modeling the complex dependencies among chemical species. To fill this gap, we propose a Bayesian causal regression factor model that estimates the multivariate causal effects of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States. Our approach introduces two key innovations: (i) a causal inference framework for multivariate potential outcomes, and (ii) a novel Bayesian factor model that employs a probit stick-breaking process as prior for treatment-specific factor scores. By focusing on factor scores, our method addresses the missing data challenge common in causal inference and enables a flexible, data-driven characterization of the latent factor structure, which is crucial to capture the complex correlation among multivariate outcomes. Through Monte Carlo simulations, we show the model’s accuracy in estimating the causal effects in multivariate outcomes and characterizing the treatment-specific latent structure. Finally, we apply our method to US air quality data, estimating the causal effect of wildfire smoke on 27 chemical species in pm2.5, providing a deeper understanding of their interdependencies.
footnotetext: dafne__\__zorzetto@brown.edu
1
Data Science Institute, Brown University, Providence, Rhode Island, U.S.A.
2 Department of Biostatistics, Harvard University, Cambridge, Massachusetts, U.S.A.
3 Department of Biostatistics, Brown University, Providence, Rhode Island, U.S.A.
4 Department of Data Science, Dana Farber Cancer Institute, Boston, Massachusetts, U.S.A.
Keywords: Causal inference, Factor analysis, Factor score’s prior, Infinite mixture distribution, Potential outcome framework.
1. Introduction
Wildfires have emerged as a critical environmental and public health concern, significantly deteriorating air quality through the emission of smoke and associated pollutants (Childs et al., 2022; Krasovich Southworth et al., 2025). In the United States, wildfire smoke exposure has increased significantly, with the burned area quadrupling over the past four decades—a trend expected to continue as climate conditions become increasingly favorable to fire activity. Wildfire smoke can rise high into the atmosphere and travel vast distances, affecting air pollution levels and health risks far from the source (Burke et al., 2021; O’Dell et al., 2021). As a results, it has received considerable attention in environmental health research (Reisen et al., 2015; Burke et al., 2021; Krasovich Southworth et al., 2025; Aguilera et al., 2021).
A growing body of literature has examined the effects of wildfire smoke on particulate matter (pm2.5), highlighting the need to quantify the proportion of pm2.5 attributable to wildfires (Burke et al., 2023; Childs et al., 2022), given the well-documented adverse health effects of elevated pm2.5 exposure (Carone et al., 2020; U.S. Environmental Protection Agency, 2022). However, wildfire smoke is a complex mixture of pollutants that undergoes physical and chemical transformations during atmospheric transport (Reisen et al., 2015). Understanding how wildfire smoke influences the concentration levels of various chemical species in pm2.5is therefore a pressing challenge.
While previous studies have focused on the impact of wildfire smoke on individual chemical components—such as phosphorus and nitrogen (Spencer and Hauer, 1991), lead (Odigie and Flegal, 2014), bromine and other metals (Li et al., 2023)—the recent work of Krasovich Southworth et al. (2025) has provided a more comprehensive analysis by quantifying the contribution of wildfire smoke to the concentration of 27 different chemical species in pm2.5 using a standard linear regression model. Their analysis primarily focuses on accounting for temporal variation and estimating the contribution of wildfire smoke to each chemical species in pm2.5. However, it does not jointly model the chemical species, limiting information sharing across species, and it is not adopting a causal inference prospective, restricting its ability to disentangle the causal effect of wildfire smoke.
A causal inference framework is crucially needed to rigorously assess and quantify the effect of wildfire smoke on changes in chemical concentration levels, providing a deeper understanding of the associations highlighted in previous studies. Moreover, incorporating the high correlation structure among the different chemical species is critical to capturing deeper insights of the smoke composition and its broader environmental effects.
To address these limitations, we propose a novel Bayesian factor model for causal inference, which leverages infinite mixture distributions as priors for the treatment-specific factor scores. The key novelty of our project lies in two main aspects. First, we define a causal inference framework in a Bayesian factor analysis model to account for the high correlation among multivariate potential outcomes, such as the chemical species in pm2.5. Second, we introduce a novel prior on the factor scores using a Dirichlet process (Quintana et al., 2022) with a probit linear model to capture treatment-specific heterogeneity in a data-driven way.
Traditional Bayesian factor models have focused primarily on priors for factor loadings (Bhattacharya and Dunson, 2011; Schiavon et al., 2022; Frühwirth-Schnatter et al., 2024; De Vito et al., 2021), typically assuming standard normal distributions for factor scores. Recent works (Zorzetto et al., 2024; Bortolato and Canale, 2024) have begun exploring priors on factor scores, since they define the quantity that each subject has on the coresponding factor, making them essential in many applications. However, the causal inference framework in Bayesian factor models introduces new challenges due to the fundamental missing data problem of causal inference (Rubin, 1986). Our proposed prior addresses this by allowing covariate-dependent structure in treatment-specific scores, enabling better estimation of latent confounding patterns.
The paper is organized as follows. Section 2 introduces the causal inference setup, including the estimands of interest, causal assumptions, and identification strategy. Section 3 introduces our Bayesian factor model for causal inference, highlighting the innovation in our prior specification for factor scores and its critical role in this context. In Section 4, we assess the model’s performance through simulations, comparing it to state-of-art approaches in causal inference and evaluating the importance of our proposed prior for the factor score compared to the factor model with the standard prior. In Section 5, we apply our approach to estimate the causal effect of the presence of wildfire smoke on the concentration of 27 chemical species in pm2.5 across the United States, using data from common sources of Krasovich Southworth et al. (2025): F.E.D. IMPROVE (2024) and Childs et al. (2022). We conclude the paper with a discussion and future directions in Section 6.
2. Causal setup
Let $i$ denote the study unit, with $i=1,\dots,n$, and $T_{i}\in\{0,1\}$ the binary treatment random variable with observed value $t_{i}$. Let also ${\bf X}_{i}=(X_{1i},\dots,X_{pi})^{T}$ denote a vector of $p$ observed covariates (confounders), which may be either continuous or discrete, and $\mathbf{Y}_{i}\in\mathbb{R}^{q}$ the multivariate outcome of dimension $q$, where the outcome components are potentially highly correlated, a structure exploited by our Bayesian causal factor model. Throughout the paper, bold letters denote vectors.
Following the definition of the Rubin Causal Model (Rubin, 1974), we invoke the Stable Unit Treatment Value Assumption (SUTVA; Rubin, 1980), which is a combination of two assumptions: no interference among the units, i.e. the potential outcome values from unit $i$ do not depend on the treatment applied to other units, and consistency, i.e. each observed outcome is the individual’s potential outcome under their treatment. Under SUTVA, we define the potential outcomes, for unit $i$, as $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, where ${\bf Y}_{i}(t)=(Y_{i1}(t),\dots,Y_{iq}(t))^{T}$ for $t\in\{0,1\}$. Specifically, ${\bf Y}_{i}(0)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the control group and ${\bf Y}_{i}(1)\in{\mathds{R}}^{q}$ is the $q$-variate outcome when the unit $i$ is assigned to the treatment group. Under consistency, therefore the observed outcome can be written as:
${\bf Y}_{i}=(1-T_{i})\,{\bf Y}_{i}(0)+T_{i}\,{\bf Y}_{i}(1).$
In practice, for $i=1,\dots,n$, we observe only ${\bf y}_{i}\in{\mathds{R}}^{q}$, the realization of the random variable ${\bf Y}_{i}=(Y_{i1},\dots,Y_{iq})$. Conversely, we do not observe the missing outcome ${\bf Y}_{i}^{mis}\in{\mathds{R}}^{q}$ defined as ${\bf Y}_{i}^{mis}=T_{i}\,{\bf Y}_{i}(0)+(1-T_{i})\,{\bf Y}_{i}(1)$.
Moreover, we assume an unmeasured variable ${U}_{i}$ for each unit $i$, which affects the treatment $T_{i}$ and the potential outcome $\{{\bf Y}_{i}(0),{\bf Y}_{i}(1)\}$, the last exclusively through latent factors $\{{\bf L}_{i0},{\bf L}_{i1}\}$. In particular, Figure 1 illustrates the three cases considered in our causal setup.
Several key clarifications are necessary, with a more detailed explanation provided in the following sections. First, although we refer to $U$ as a single random variable, it can be generalized to a function of multiple unmeasured variables. Second, the latent factors serve as ‘proxy’ variable that adjust the unmeasured variable, leveraging the correlation among the multivariate outcomes. Third, any relations between the unmeasured variable ${U}_{i}$ and the confounders ${\bf X}_{i}$ are considered, as reported in Figure 1.
Therefore, when $U$ affects the outcome only via the latent factors—as assumed in this paper—then this unmeasured information can be recovered through the latent structure estimated in our proposed Bayesian causal factor model, formally introduced in Section 3.
Figure 1. Graphical representation of the causal pathway assumed in this paper with treatment $T$, measured confounders $X$, unmeasured confounders $U$, latent factors $L$, and multivariate outcome $Y$. The three scenarios consider different relationships of unmeasured confounders $U$: (1) confounding only through an effect on latent factors $L$, (2) confounding as an effect of measured confounders $X$, (3) cause of measured confounders $X$. The simulation study investigates each of these scenarios.
Causal Estimands and Identifying Assumptions
Our goal is to estimate the causal effect of the treatment on each element of the multivariate outcome. Therefore, we define the quantity of interest as the Sample Average Treatment Effects (SATE)
$\mbox{{SATE}}={\mathds{E}}[{\bf Y}(1)-{\bf Y}(0)],$
where the average is taken across the units $\{1,\dots,n\}$ and SATE $\in\mathbb{R}^{q}$ is the vector of causal effects for the components of the outcome.
As common in causal inference, these causal estimands cannot be estimated directly from observed data due to the missing potential outcomes. To address this, we adopt the following assumptions and leverage the features of our proposed factor model, specifically the relationship of treatment-specific factor scores—i.e., ${\bf l}_{0}$ and ${\bf l}_{1}$—with measured and unmeasured variables. The first two assumptions are taken from standard causal inference literature, and the third is newly defined in the context of the assumed latent factor structure.
Positivity assumption. Each unit $i\in\{1,\dots,n\}$ has a non-zero probability o of receiving each treatment level,
$10<\Pr(T_{i}=1\mid{\bf X}_{i}={\bf x}_{i})<$ 1 .
Conditional ignorability assumption. The assignment of treatment is random in each group of units characterized by observed covariates ${\bf x}$ and latent factors scores ${\bf l}_{0}$ and ${\bf l}_{1}$, such that for each unit $i\in\{1,\dots,n\}$
$\{{\bf Y}_{i}(1),{\bf Y}_{i}(0)\}\perp T_{i}\mid\{{\bf X}_{i}={\bf x}_{i},{\bf L}_{i0}={\bf l}_{i0},{\bf L}_{i1}={\bf l}_{i1}\}.$
Indirect unmeasured confounding assumption. The unmeasured variable $U$ affects the potential outcomes only indirectly, through the latent factor and the observed confounders, such as
$\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1},U)=\mathds{P}({\bf Y}(1),{\bf Y}(0)\mid{\bf X},{\bf L}_{0},{\bf L}_{1})$
This assumption can be visualize in Figure 1 from the absence of a direct arrow between the $U$ and ${\bf Y}$.
Identification
Under the three assumptions from previous Subsection, the causal estimands (i.e., functions of potential outcomes) can be re-expressed as statistical estimands (i.e., functions of random variables) and can then be estimated using the observed data.
Property 1.
Invoking the previous assumptions, we can obtain the statistical estimand for SATEs as the following
${\mathds{E}}[{\bf Y}(1)-{\bf Y}(0)]$
$\quad=\int_{\mathcal{X}}\int_{\mathbb{R}^{J_{0}}}\int_{\mathbb{R}^{J_{1}}}\bigg{(}{\mathds{E}}[{\bf Y}\mid T=1,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$
$\quad\quad\quad\quad-{\mathds{E}}[{\bf Y}\mid T=0,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]\bigg{)}$
×$\quad\quad\quad\quad\times\mathds{P}({\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}\mid{\bf X}={\bf x})\mathds{P}({\bf X}={\bf x})d{\bf l}_{0},d{\bf l}_{1}d_{\bf x};$
where the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$, for each treatment level $t\in\{0,1\}$, will be estimated using our proposed factor model (2) and the joint probability of the latent factor score $\mathds{P}({\bf L}_{0},{\bf L}_{1}\mid{\bf X})$ modeled via prior distribution specified in (3).
A similar expression applies for the SATEk; see Appendix B for further details and full proof of Property 1.
Property 1 underlines the pivotal role of the probability distribution of treatment-specific factor scores, and thus the necessity of defining a flexible prior, such as the one we propose in in (3). Moreover, the inner expectation ${\mathds{E}}[{\bf Y}\mid T=t,{\bf X}={\bf x},{\bf L}_{0}={\bf l}_{0},{\bf L}_{1}={\bf l}_{1}]$ can be rewritten as conditional on the unmeasured variable $U$, as follows:
${\mathds{E}}[{\bf Y}\mid t,x,l_{0},l_{1}]$
[…]
where the indirect unmeasured confounding assumption assumption is invoked for the second step. The involvement of the unmeasured variable $U$ in the conditional probability of factor scores $\mathds{P}(l_{0},l_{1}\mid u,x)$ leads to a careful choice of the distribution of this probability. We postpone the discussion to Section 3.
3. Bayesian Causal Factor Model
Model definition
Due to the multivariate nature of the outcome, we propose a factor analysis based approach (Thurstone, 1931; Anderson-Cook, 2004) to model the potential outcomes given the treatment and the confounders. Rather than performing multiple univariate analyses, our factor analysis approach (i) improves statistical power by leveraging the correlation structure among the multivariate outcome, and (ii) facilitates the recovery of information from the unmeasured variable ${U}$.
Specifically, for each unit $i\in\{1,\dots,n\}$, the outcome vector ${\bf Y}_{i}$ follows a multivariate normal distribution such that
$\{{\bf Y}_{i}\mid{\bf x}_{i},t,{\bf l}_{it}\}\sim g_{t}({\bf x}_{i},{\bf l}_{it}),$
(2) $g_{t}({\bf x}_{i},{\bf l}_{it})=\boldsymbol{\mu}_{t}+{\bf B}_{t}{\bf x}_{i}+\Lambda_{t}{\bf l}_{it}+\boldsymbol{\xi}_{it},$
$\boldsymbol{\xi}_{it}\sim\mathcal{N}_{q}(0,\Psi_{t}).$
The outcome ${\bf Y}_{i}$ depends on treatment $t$, covariates ${\bf x}_{i}$, and treatment-specific latent factors ${\bf l}_{it}\in\mathbb{R}^{J_{t}}$, where $J_{t}$ denotes the number of factors for treatment $t$ (assumed unknown). The matrix $\Lambda_{t}\in\mathbb{R}^{q\times J_{t}}$ represents treatment-specific factor loadings, and $\Psi_{t}\in\mathbb{R}^{q\times q}$ is assumed to be a diagonal matrix with elements $(\psi_{1t},\dots,\psi_{qt})$ as in standard factor analysis literature (Thurstone, 1931; Anderson-Cook, 2004).
[…]
$\boldsymbol{\mu}_{t}\sim\mathcal{N}_{q}(\boldsymbol{\mu}_{m_{t}},\Sigma_{m_{t}})\mbox{ and }\boldsymbol{\beta}_{tj}\sim\mathcal{N}_{q}(\boldsymbol{\mu}_{\beta_{t}},\Sigma_{\beta_{t}}),$
where $\boldsymbol{\beta}_{tj}$ denotes the $j$-th row of the matrix ${\bf B}_{t}$ in (2), such that $\boldsymbol{\beta}_{tj}=\{\beta_{tj1},\dots,\beta_{tjp}\}$ and $\beta_{tjk}$ is the element $({\bf B}_{t})_{j,k}$ with $k\in\{1,\dots,p\}$. We consider independent priors such that $\Sigma_{m_{t}}$ and $\Sigma_{\beta_{t}}$ are diagonal matrices for each $t$. Alternative priors—e.g., mixtures of Gaussian distributions or spike and local prior—could be employed.
The Role of the DDP in the Factor Score
In the literature, Bayesian statistical inference in factor analysis primarily focuses on prior for factor loading matrix, whereas the prior for the factor score matrix is typically non-informative and chosen as a standard Gaussian distribution, assuming independence across elements.
However, this prior is unsuitable for handling missing data, as required in the potential outcomes framework for causal inference. Since accurate imputation of missing data is crucial, an informed choice of prior distribution is necessary. To address this, we propose a prior for the treatment-specific factor scores that takes advantage of the fundamental properties of the dependent Dirichlet process (DDP; Quintana et al., 2022). In particular, the DDPs are particular infinite mixture models—see details in the following Subsection—where our specific choice is to incorporate covariates into the weights to enhance the precision of imputation.
This capability is crucial, as it allows for the modeling of complex probability distributions while leveraging covariates to fully characterize them. This distinct advantage has made DDPs a widely adopted and powerful tool in causal inference settings involving scalar outcomes. For example, Zorzetto et al. (2024) introduces the confounder-dependent Bayesian mixture model where the DPSB identifies groups in the population that characterize the heterogeneity in causal effects, the enriched Dirichlet process is used by Roy et al. (2018) and Roy et al. (2024), respectively, to use DDPs for missing in random covariates and mediation analysis, and Schwartz et al. (2011) and Zorzetto et al. (2024) specify a DDP prior for principal stratification.
In our factor model, we want to specify a DDP prior that allows us to capture heterogeneity at both the factor and treatment levels. The observed data inform the heterogeneity in the probability distribution of factor scores, facilitating the identification and estimation of data-driven clusters. Specifically, covariates determine the clustering of units through the mixture weights (see Eq. (5)), ensuring that units with similar covariate profiles share cluster-specific characteristics relevant for imputing missing potential outcomes.
This data imputation process can be viewed as analogous to the matching process typically used in study design, where the criteria are driven by the covariates. However, our prior specification (Eq. (3)) does not replace the study design in observational studies, as further discussed in the application section.
Depend Mixture Distribution for Treatment-Specific Factor Scores
We define the prior for each factor $l_{itj}$ as follows
$\{l_{itj}\mid{\bf x}_{i},t\}\sim h_{tj}(\cdot\mid{\bf x}_{i}),$
(3) $h_{tj}(\cdot\mid{\bf x}_{i})=\int_{\Omega}{\mathcal{H}}(\cdot;\omega)dG_{{\bf x}_{i}}^{(tj)}(\omega),$
$G_{{\bf x}_{i}}^{(tj)}\sim\Pi_{{\bf x}_{i}}^{(tj)},$
for $j\in\{1,\dots,J_{t}\}$ of units $i\in 1,\dots.n\}$, for each $t\in\{0,1\}$. The random probability measure $G_{{\bf x}_{i}}^{(tj)}$ depends on confounders ${\bf x}_{i}$ and is specific to a treatment level $t$ and a factor $j$. Assuming its prior $\Pi_{{\bf x}_{i}}^{(tj)}$ where $\Pi_{{\bf x}_{i}}^{(tj)}$ is a DDP (Quintana et al., 2022), we define a dependent infinite mixture distribution as the probability density distribution of the factors.
To align with the standard factor analysis literature, we define the continuous density function ${\mathcal{H}}(\cdot;\omega)$, with parameter space $\Omega$, as a Gaussian kernel, where $\omega=(\eta,1/\tau)$ are the position and scale parameters, respectively.
Following a single-atom DDP (Quintana et al., 2022) characterization of the random measure $G_{{\bf x}_{i}}^{(tj)}$, we can write:
(4) $G_{{\bf x}_{i}}^{(tj)}=\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\delta_{\mathbb{\omega}^{(tj)}_{r}},$
for $t=\{0,1\}$ and $j=\{1,\dots,J_{t}\}$, where $\{\pi_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ and $\{\omega^{(tj)}_{r}\}_{l\geq 1}$ represent infinite sequences of random weights and random kernels’ parameters, respectively. Notably, both random sequences depend on the level of treatment $t$ and the factor $j$, while the weights also depend on the values of the confounders ${\bf x}_{i}$.
The sequence of random parameters $\{\omega_{r}^{(tj)}\}_{r\geq 1}$ is independent and identically distributed, where each $r$-th element of the sequence is the joint parameters of the $r$-th Gaussian distribution in the mixture model in (3), such that $\omega_{r}^{(tj)}=(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)})$. Following the stick-breaking representation (Sethuraman, 1994), the sequence of dependent weights is defined as
$\pi_{r}^{(tj)}({\bf x}_{i})=V^{(tj)}_{r}({\bf x}_{i})\prod_{g<r}\{1-V_{g}^{(tj)}({\bf x}_{i})\},$
where $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ are $\{0,1\}$-valued independent stochastic processes.
Among the dependent nonparametric processes, the Dependent Probit Stick-Breaking (DPSB) (Rodriguez and Dunson, 2011) defines the process $\{V_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ as follows
$V_{r}^{(tj)}({\bf x}_{i})=\Phi(a_{r}^{(tj)}({\bf x}_{i})),$
(5) $a_{r}^{(tj)}({\bf x}_{i})\sim\mathcal{N}(\alpha_{0r}^{(tj)}+\alpha_{1r}^{(tj)}{\bf x}_{1i}+\dots+\alpha_{pr}^{(tj)}{\bf x}_{pi},1),$
where $\Phi(\cdot)$ is the Probit function and $\{a_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates ${\bf x}_{i}$.
The nature of the prior distribution for the treatment-specific factors (3) allows us to introduce a latent categorical variable $S_{itj}$, for each factor $j\in\{1,\dots,J_{t}\}$ treatment $t\in\{0,1\}$ unit $i\in\{1,\dots.n\}$, describing clusters of units defined by heterogeneous responses to the treatment level $t$. Assuming $\Pr(S_{itj}=r)=\pi_{r}^{(tj)}({\bf x}_{i})$, then the prior distribution of the treatment-specific factors (3) can be rewritten, conditional on $S_{itj}$, as
(6) $\{{\bf l}_{itj}\mid S_{itj}=r,\boldsymbol{\omega}\}\sim\mathcal{N}(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)}).$
Statistical Guarantees
The proposed prior for treatment-specific factor scores offers several strengths, as outlined in the previous Subsection. Additionally, it serves as a powerful tool for inferring information about the relationships among the multiple outcomes, as governed by unobserved features $U$. Specifically, the flexibility of the DDP distribution enables it to capture this unmeasured information while effectively controlling bias.
In particular, we consider the case where $U\sim\mathcal{N}(\mu_{u},\sigma_{u}^{2})$ with non-negative correlation $\sigma^{(lu)}$ between $U_{i}$ and $l_{itj}$, allowing us to derive closed-form expressions that clarify the relationships among the quantities involved.
Indeed, using multivariate normal properties, we can express the marginal distribution of treatment-specific factor scores ${\bf l}_{itj}$ and $U_{i}$, for each unit $i$, treatment $t$, and factor $j\in\{1,\dots,J_{t}\}$, so we can rewrite the marginal distribution of treatment-specific factor scores ${\bf l}_{itj}$ and $U_{i}$, for each unit $i$, treatment $t$, and factor $j\in\{1,\dots,J_{t}\}$, as follows:
(11) $\left[\begin{array}[]{c}{\bf l}_{itj}\mid{\bf x}_{i}\\ U_{i}\end{array}\right]\sim\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\mathcal{N}\left(\left[\begin{array}[]{c}\eta_{r}^{(tj)}\\ \mu_{u}\end{array}\right],\Sigma_{r}^{(LU)}\right),$
(14) $\Sigma_{r}^{(LU)}=\left[\begin{array}[]{cc}1/\tau_{r}^{(tj)}&\sigma^{(lu)}\\ \sigma^{(lu)}&\sigma_{u}^{2}\end{array}\right].$
By the the properties of the multivariate normal distribution, the conditional distribution of the factor scores given the covariates and the unmeasured variable is given by
(15) $\{{\bf l}_{itj}\mid{\bf x}_{i},U_{i}\}\sim\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\mathcal{N}(\gamma_{0r}+\gamma_{1r}U_{i},\sigma^{(l|u)2}_{r}),$
where the location and scale parameters are defined as following
$\gamma_{0r}+\gamma_{1r}U_{i}=\eta_{r}^{(tj)}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}(U_{i}-\mu_{u}),$
$\sigma^{(l|u)2}_{r}=\frac{1}{\tau_{r}^{(tj)}}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}.$
Although the conditional distribution in Equation (15) cannot be directly estimated due to the unmeasured variable due to the unmeasured variable $U$, the following property 2 demonstrates that our proposed formulation for the distribution of treatment-specific factors ${\bf L}_{t}$ is sufficient to recover the relevant information about $U$.
Property 2.
Assuming the unmeasured variable $U$ is normal distributed, the treatment-specific factors ${\bf L}_{t}$, for each $t$, are defined as in (3)-(5), and the assumptions of Eq.(11) hold, then each component of the mixture $r\geq 1$, has: (i) a finite expected value of the location parameter, that takes into account the expected value of the unmeasured variable $U$, and (ii) the scale parameter has the scale parameter of the conditional distribution (15) as a lower bound. Specifically,
$(i)\quad{\mathds{E}}[\eta_{k}]=\gamma_{0r}+\gamma_{1r}\mu_{u}={\mathds{E}}[{\bf l}_{tj}\mid{\bf x},U],$
(16) $(ii)\quad\frac{1}{\tau_{r}^{(tj)}}=\sigma^{(l|u)2}_{r}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}\geq\sigma^{(l|u)2}_{r}.$
Therefore, the probability distribution of $\{L_{tj}\mid{\bf X}\}$ has same mean of $\{L_{tj}\mid{\bf X},U\}$ and an equal or larger variance. Substituting this result into Eq. (1), we conclude that the casual effects—the SATE—are unbiased under the assumptions defined in Section 2.
Scenario 1 in Figure 1 corresponds to the case where factor scores and the unmeasured variable are independent, i.e., $\sigma^{(lu)}=0$. In this case, the marginal and conditional variances are the same, indeed the equation (16) can be simplified in $1/\tau_{r}^{(tj)}=\sigma^{(l|u)2}_{r}$. Conversely, Scenarios 2 and 3 assume $\sigma^{(lu)}>0$. Assuming a positive value for $\sigma^{(lu)}$ does not determinate the direction of dependence, thus encompassing both cases.
Shrinkage Priors for Factor Loading
For the factor loading matrix, we adopt the well-known shrinkage prior introduced by Bhattacharya and Dunson (2011). For each treatment-specific factor loading $\lambda_{tjh}$, where treatment level $t\in\{0,1\}$, the factor $j=1,\ldots,\infty$ and the outcome variable $h\in\{1\ldots p\}$:
$\lambda_{tjh}\mid k_{tjh},\iota_{th}\sim N\left(0,\theta_{tjh}^{-1}\iota_{th}^{-1}\right),$
$\theta_{tjh}\sim\operatorname{Gamma}\left(\nu_{t}/2,\nu_{t}/2\right),$
$\iota_{th}=\prod_{l=1}^{h}\delta_{tl},$
$\delta_{t1}\sim\operatorname{Gamma}\left(a_{1_{t}},1\right),\;\delta_{tl}\sim\operatorname{Gamma}\left(a_{2_{t}},1\right)\forall l\geq 2.$
Alternative priors include the Bayesian cumulative shrinkage introduced by Legramanti et al. (2020) or a generalization of Bhattacharya and Dunson (2011)’s prior proposed by Schiavon et al. (2022).
4. Simulation Study
We evaluate the performance of the proposed model through an extensive simulation study. The primary objective is to assess its ability to estimate causal effects, with a particular focus on controlling the variability of SATE and adjusting the bias induced by the unmeasured variable $U$. We compare our causal factor model against state-of-the-art flexible models in causal inference—causal BART by Hill (2011) and Bayesian causal forest (BCF) introduced by Hahn et al. (2020)—as well as a standard factor model employing a standard normal prior for factor scores and adjusted for the causal inference setting. Specifically, the standard factor model is defined as in (2), but with priors ${\bf l}_{it}\sim\mathcal{N}(0,1)$ for each unit $i$ and treatment level $t\in\{0,1\}$. This setup allows us to compare and evaluate the impact of the proposed DDP prior on treatment-specific factor scores. While BART and BCF are estimated separately for each element of the outcome variable, both the standard factor model adjusted for causal inference and our proposed approach explicitly account for the multivariate nature of the outcome.
We consider four distinct simulation scenarios. The first three explore different relationships between the unmeasured variable $U$ and the confounders ${\bf X}$, corresponding to the graphical representations in Figure 1. The fourth scenario is designed to closely mimic the real dataset used in our application in Section 5. Details of simulation-generating processes are provided in Table 1.
Table 1. Parameters and distributions for the simulation scenarios.
Scenario 1 Scenario 2 Scenario 3 Scenario 4
Dimensions:
$n$ 500 500 500 3426
$J_{t}$ $(3,3)$ $(3,3)$ $(3,3)$ $(3,3)$
$q$ 10 10 10 27
$p$ 4 4 4 28
Variables:
$U$ $U\sim N(0,2)$ $U\sim N(f_{U}(X_{1:4}),0.5)$ $U\sim N(0,2)$
${\bf X}$ $X_{1:2}\sim N(0,1)$ $X_{1:2}\sim N(0,1)$ $X_{k}\sim N(f_{k}(U),1),\,k=1,2$ observed ${\bf X}$
$X_{k}\sim Be(\pi_{k}),\,k=3,4$ $X_{k}\sim Be(\pi_{k}),\,k=3,4$ $X_{k}\sim Be(\pi_{k}),\,k=3,4$
$T$ $T\sim Be(f_{T}(X_{1:4},U))$ $T\sim Be(f_{T}(X_{1:4},U))$ $T\sim Be(f_{T}(X_{1:4},U))$ observed $T$
Factors:
Clusters 3 or 2 cluster for each treatment-specific factors: $C_{ith}:=f_{C}({\bf X}_{1:4})$ $C_{ith}:=f_{C}({\bf X}_{1:27})$
$l_{it}$$\{l_{ith}\mid C_{ith}=c\}\sim N(\mu_{tc}+\gamma_{tc}U,1),\;h=\{1,\dots,p\}$$\{l_{itj}\mid C_{ith}=c\}\sim N(\mu_{tc},1)$
$\Lambda_{t}$ matrices with $25\%$ of zeros and nonzeros elements generated estimated $\Lambda_{t}$
by $\lambda_{th}\sim\pi Unif[-1,-0.8]+(1-\pi)Unif[0.8,1]$, with $\pi\sim Be(0.5)$ in observed data
Outcome mode:
$B_{t}$ $\beta_{ht}\sim Unif[-3+t,2+t],\;t=0,1,\;h=1,\dots,p$
$\epsilon_{t}$ $\xi_{it}\sim Unif_{p}[0,1],\;t=0,1$
${\bf Y}$ $Y_{i}(t)=B_{t}X+\Lambda_{t}l_{it}+\xi_{it},\;t=0,1\;i=1,\dots,n$
Specifically, in the first three scenarios, the treatment-specific factor scores $\{{\bf L}_{0},{\bf L}_{1}\}$ depend on both the confounders ${\bf X}$ and the unmeasured variable $U$, while the potential outcomes ${\bf Y}$ depend on $U$ only through the factors. In Scenario 1, illustrated on the left of Figure 1, ${\bf X}$ and $U$ are independent (i.e., $\sigma^{lu}=0$ in Property 2). Scenario 2 assumes $U$ depends on ${\bf X}$, and Scenario 3 explores the reverse relationship, where the confounders ${\bf X}$ depend on $U$. These dependencies are highlighted by the red dotted arrows in the second and third representations of Figure 1. To simulate that, as reported in Table 1, Scenario 2 has four independent confounders ${\bf X}$ and $U$ dependent on them, while Scenario 3 is defined with an independent normal distribution for $U$ and the first two confounders ${\bf X}$ as normal distributed with mean regression dependent of $U$. Scenario 4 uses real-data information for the 28 confounders and the treatment assignment. Factor loadings are estimates with our model applied to real data. Factor scores and outcomes were simulated according to our proposed model.
For each of the four scenarios, we generated 50 collections of datasets. Causal effects are estimated using the R package bartCause to estimate with the BART model, the bcf code available on GitHub for BCF estimation, and our own Gibbs sampler for the factor models, available on GitHub at dafzorzetto/BayesCausalFactor. The same codebase includes the implementation of the standard factor model.
We evaluate the performance and accuracy of the models according to the bias mean and square error (MSE) of SATEk  estimation for each $k=\{1,\dots,10\}$ in Scenarios 1–3, and $k=\{1,\dots,27\}$ in Scenario 4. The results are summarized in Figure 2. Our proposed model (CausalFA) outperforms the competitors in both bias and MSE, showing unbias estimations of the causal effects and smaller variability. Specifically, BART and BCF exhibit an overestimation of variability, leading to high MSE values, likely due to their independent estimation of each causal effect. The factor model with the standard normal prior (StandardFA) reveals bias estimation in the causal effects, highlighting a critical weakness of the standard prior for the factor scores and the benefit of the DDP prior.
Our proposed model is the only one among the four implemented models to provide an unbiased estimation of the causal effects for all the elements in the multivariate outcomes across the simulated scenarios, demonstrating its ability to adjust the estimation for the unmeasured variable through the latent factors and their prior.
Figure 2. Bias (left) and mean square error (MSE, right) across simulated scenarios. Results are shown for our proposed model, standard factor model, causal BART, and BCF.
5. Environmental Application
Wildfires have become a major environmental and public health concern, significantly deteriorating air quality through the release of smoke and associated pollutants. Although previous studies have examined the broad effects of wildfire smoke on particulate matter (pm2.5), a detailed understanding of its causal effect on specific chemical species in the environment remains limited. The recent work of Krasovich Southworth et al. (2025) laid the foundation for studying wildfire smoke and chemical variation in the United States, highlighting strong interconnections between them. However, their analysis is based on temporal correlations adopting a linear modeling approach for each individual chemical. Our study addresses two distinct objectives employing a novel methodology. First, we aim to answer the causal question: What is the causal effect of wildfire smoke on each of the chemicals on pm2.5? Second, we explicitly account for the correlation between chemicals, providing robust estimations through our proposed causal factor model.
Following Krasovich Southworth et al. (2025), who report strong temporal correlations in the data, we focus on a specific time period—specifically, the summer months—when wildfires are more frequent. Moreover, in their analysis they highlight the difficulties in defining the level of pm2.5 induced by wildfire smoke and of other origins due to different measurement technologies. To account for this, our treatment variable is defined as the binary presence or absence of wildfire smoke.
Data
The data analyzed are merged from different sources, following a strategy similar to Krasovich Southworth et al. (2025). We consider the $280$ air pollution monitors across the United States, operated by the National Park Service and the U.S. Forest Service. The data are collected from the Environmental Protection Agency’s (EPA) pm2.5 Chemical Speciation Network (CSN) and the Interagency Monitoring of Protected Visual Environments (IMPROVE) program, and stored in the Federal Land Manager Environmental Database (F.E.D. IMPROVE, 2024). We select the same $27$ chemical species considered in Krasovich Southworth et al. (2025), divided in alkaline-earth metals (Magnesium, Calcium, and Strontium), alkali metals (Sodium, Rubidium, and Potassium), transition metals (Chromium, Nickel, Vanadium, Copper, Iron, Zinc, and Manganese), metalloids (Arsenic and Silicon), other metals (Lead, Aluminum, and Titanium), nonmetals (Selenium, Nitrate, Sulfate, Sulfur, and Phosphorus), halogens (Chlorine and Bromine), and Organics (elemental Carbon and organic Carbon).
Each monitor records daily information on the level of each chemical species, which is collected every three days. The IMPROVE and CSN monitoring data were preprocessed to replace data flagged as unacceptable quality with interpolated values or, when appropriate, a value equal to 1/2 of the minimum detection limit (less than 3% of values were interpolated). To adjust for methodological differences between the CSN and IMPROVE monitors, we include the variable monitor_type as a confounder.
Wildfire smoke pm2.5 concentration are derived from the daily prediction of Childs et al. (2022), available from the GitHub repository echolab-stanford/daily-10km-smokePM. We define the treatment as $T=1$ for “smoke days” with non-zero predictions and control $T=0$ for “non-smoke days” with zero predicted wildfire smoke pm2.5. The prediction resolution is in the grid $10Km$ (i.e., each unit aggregates the information over $10Km\times 10Km$ area), therefore we match the location of each air pollution monitor with the closest point of the grid.
As confounders, we collect weather information (from the Harvard Dataverse (Audirac, 2024)) for U.S. ZIP Code Tabulation Areas (ZCTA) at daily resolution and census data (via the tidycensus R package (Walker et al., 2021)). Weather variables include daily values for near-surface air temperature, maximum near-surface air temperature, precipitation, minimum and maximum near-surface relative humidity, surface downwelling solar radiation, and wind speed and direction at $10m$. Census variables include population, percentage of males, percentage of different ethnicities (White, Black and African American, Asian, and other), median household income, total number of housing units, and poverty status. Both the weather information and census data are matched with the monitor locations.
Study design
Our analysis focuses on July–September 2014, when wildfire smoke is most prevalent in the United States (Krasovich Southworth et al., 2025). The observational unit is a monitor-day. Since chemical concentrations are measured every three days, the dataset contains $7,467$ units.
The treatment variable is defined as wildfire smoke exposure $T=0$ for non-smoke days and $T=1$ for smoke days. The multivariate outcome is the concentration of $27$ chemical species, measured in $\mu g/m^{3}$ and transformed on the natural logarithmic scale. Confounders include weather information, census data, type of monitor indication, month, latitude and longitude of monitor location, and type of monitor location (rural, city or intermediate).
Initially we have $5,754$ monitor-day units assigned in the control group—that is, no wildfire smoke day in that monitor location—and $1,713$ in the treated group—that is, exposed to wildfire smoke. However, we use matching before running our model to make our analyses as robust as possible with respect to the potential measured confounding bias, using confounders to measure the similarity of the units in the two groups. This is common in observational studies (Rosenbaum and Rubin, 1983) and in research on air pollution effects on health (see, e.g., Lee et al., 2021; Zorzetto et al., 2024).
We employ a 1-to-1 nearest neighbor propensity score matching, yielding a balanced sample of $3,426$ units, and improving the balance of the covariates. The reduction in units is due to the different sample sizes of the treated and control groups in the original data, and 1-to-1 matching creates a sample with the same size for the treated and control groups. The causal effects, estimated in the following section, target the sample obtained through the matching procedure.
Results
Similarly to the simulation study, we estimate the causal effect of wildfire smoke on $27$ chemical species using four models: our proposed causal factor model, the factor model with a standard Gaussian distribution for the factor score, BART, and BCF. The results are compared in Figure 3.
All four models consistently identify a positive causal effect of wildfire smoke on zinc, bromine, and the two organic elements—elemental and organic carbon—indicating that wildfire smoke increases the levels of these chemicals. This finding aligns with previous studies: Young and Jan (1977) and Odigie and Flegal (2014) observed increased zinc concentrations following wildfire smoke in California, Liu et al. (2014) reported elevated carbon levels and their negative impact on ecosystems, and Li et al. (2023) found higher bromine concentrations during wildfires in urban areas. However, the models do not fully agree on the detection of significant negative causal effects, i.e. a reduction in chemical concentrations due to wildfire smoke, particularly for strontium, sodium, nickel, titanium, silicon and aluminum. This discrepancy arises from the wide credible intervals produced by BART, which often include zero, and, in some cases, by BCF as well, limiting detection of significant effects. As shown in the simulation study, these two models tend to overestimate data variability, particularly in highly correlated outcomes. Similarly for both positive effects, only our proposed model has credible intervals that do not include the zero for cooper, lead, sulfate, and phosphorum. This is consistent with previous studies, as Young and Jan (1977) and Odigie and Flegal (2014) observed notable increases in these metals, while Spencer and Hauer (1991) reported elevated phosphorus concentrations following wildfires.
Figure 3. Estimated causal effects from four models: our proposed causal factor model, factor model with standard Gaussian prior for factor scores, BART, and BCF. Dots show the median of causal effects; lines show the corresponding $90\%$ credible interval.
The varying behaviors identified by the four models can be better understood by examining the factor loading matrix, based on correlations among the 27 chemical species. Our Bayesian causal factor model framework enables us to analyze these relationships and assess how wildfire smoke alters them by estimating treatment-specific factors—i.e., distinct factors for the presence and absence of wildfire smoke. After applying a varimax transformation (Kaiser, 1960), we identify $3$ factors for each treatment level, explaining a total of $85\%$ and $88\%$ of the variance, respectively for wildfire smoke and non-smoke days (Figure 4). In both treatment levels, the first factor captures strong correlations among transition metals, metalloids, and other metals. Notably, titanium, silicon, and aluminum—chemicals whose concentrations are found to be strongly correlated by Factor 1—are detected to have significant wildfire smoke effects by only the factor models, highlighting their ability to include correlation across outcomes and narrow uncertainty and more robust conclusions.
Factor 2 captures strong correlations among nonmetals, while Factor 3 reflects associations among organic elements. In both cases, wildfire smoke is estimated to increase their levels. This can be attributed to wildfires burning natural materials such as forests, releasing elemental and organic carbon into the air Liu et al. (2014). Likewise, nonmetals, commonly found in soil, exhibit strong positive correlations with wildfire events (Spencer and Hauer, 1991).
Furthermore, the presence of wildfire smoke appears to enhance correlations among alkaline-earth metals, transition metals, and metalloids within Factor 1, particularly among chemicals that share the same negative causal effect. Similarly, in Factor 2, wildfire smoke strengthens correlations among nonmetals, while in Factor 3, it disrupts the correlation between organic chemicals and other elements, reinforcing their internal associations. These structural shifts underscore the utility of our model in uncovering the latent structure of treatment.
Figure 4. Treatment-specific factor loadings: (left) without wildfire smoke and (right) with wildfire smoke. Colors of the chemical species names indicate chemical groupings.
6. Discussion
This work introduces two key innovations. First, from a factor analysis perspective, we propose an infinite mixture distribution for the factor scores, addressing a critical limitation in the existing literature, which has historically focused on factor loadings while largely overlooking the role of latent scores. Second, within the causal inference framework, we develop a multivariate outcome setup that enables the sharing of latent structure across outcome components, allowing for the recovery of unmeasured variables and improving the estimation of causal effects.
By combining these two contributions, the proposed Bayesian causal regression factor model demonstrates superior performance in causal effect estimation compared to state-of-the-art methods in both the factor analysis and causal inference domains, also correcting the bias in the causal effect estimates. Specifically, it outperforms standard factor models that rely on Gaussian priors for factor scores, which fail to capture heterogeneity and impute missing data. At the same time, it outperforms established methods in causal inference such as BART and BCF, which, despite their flexibility, lack the ability to share latent structure across correlated outcome components, leading to uncertainty and exhibiting weaknesses in handling variability.
While the proposed approach can be broadly applied to various real-world datasets for studying the effects of treatments or exposures on outcomes composed of correlated elements, our application in this paper focuses on a critical environmental issue. Specifically, we aim to disentangle the causal effect of wildfire smoke on the chemical composition of pm2.5, driven by awareness of the significant role that fine particulate exposure plays in human health. Our findings not only identify the chemicals that are positively or negatively affected by wildfire smoke but also reveal significant effects that other models fail to detect, due to its ability to borrow strength across outcomes. Moreover, our method effectively identifies the latent factors that drive the behavior of chemical components across different types of metals and non-metals, capturing treatment-specific patterns across classes of metals, nonmetals, and organics. This capability enriches our understanding of how wildfire smoke alters the composition of air quality in complex and systematic ways.
Our work aims to serve as a foundation for a growing body of research that highlights the importance of factor scores—supported by recent studies such as Zorzetto et al. (2024) and Bortolato and Canale (2024)—and explores the potential of factor models to quantify treatment effects in causal inference, particularly in settings where outcomes are high-dimensional and strongly correlated.
Acknowledgements
The authors thank Ting Zhang and Lucas Henneman for providing the preprocessed chemical speciation data from CSN and IMPROVE monitors.